A simple pendulum is found to vibrate 50 times within 200s. When 1.5m of its length is reduced to a certain length, it vibrates 50 times in 175s. Find the original length of the pendulum.


1
Expert's answer
2021-01-29T05:19:47-0500

50 times within 200 s corresponds to a period of


"T=\frac{200}{50}=4\text{ s}."

The original length of the pendulum can be found from the period:


"T=2\pi\sqrt{\frac{l}{g}},\\\space\\
l=\frac{gT^2}{4\pi^2}=3.97\text{ m}."